The men's world record for swimming 1500.0 m in a long course pool (as of 2007) is 14 min 34.56 s. At this rate, how many seconds would it take to swim 0.600

miles (1 mi = 1609 m)?



1
Expert's answer
2021-02-01T04:03:12-0500

14 min 34.56s = 874.56 s

0.600 miles = 0.600*1609 = 965.4 m

1500.0 m ---------- 874.56 s

965.4 m ------------- t

t = 965.4*874.56/1500.0 = 562.87 s or 9 min 22.87s

answer: 9 min 22.87 s

------------------------------------------------------------------------------------------

Another solution:

<v> = 1500.0/874.56 = 1.71515 m/s

t = 965.4/1.71515 = 562.87 s or 9 min 22.87 s

answer: 9 min 22.87 s